# Quantum chaos and the arrow of time

### v10

@@ Appendix C Comparison of canonical and microcanonical ensembles
$S_{\text{can}}=-\sum_{i}p_{i}\log p_{i},\qquad p_{i}=e^{-\beta E_{i}}$ (40)
@@
$S_{\mu}=\log W,\qquad\text{$W=$ number of microstates within $[E_{\beta}-dE/2,E_{\beta}+dE/2]$}$ (41)
@@
Figure 10: Left panels: Comparison of canonical entropy $S_{\text{can}}$ and microcanonical entropy $S_{\mu}$ as a function of inverse temperature $\beta$ for different system sizes of chaotic spin chain using different values of the energy window $dE$. Right panels: Plots of $S_{\text{can}}/S_{\mu}$ as a function of inverse temperature $\beta$ for different values of $dE$ to study $dE$-dependence of the microcanonical entropy and to find the optimum value of $dE\sim 0.8$.
Figure 11: Finding the value of the energy window $dE$ following [13]. These plots are for the system size $L=16,N=8$. Left panel: Plot of microcanonical expectation values of the kinetic energy operator using two different sizes of the energy window $dE=0.4$ and $dE=0.8$. Inset is the Log-Linear plot of the expectation value of the left ($[\bar{E}-dE/2,\bar{E}]$) and right $[\bar{E},\bar{E}+dE/2]$ averages as a function of $dE$, where $\bar{E}=-1.05$ which has effective temperature $T=3$. Right panel: Linear-Linear plot of the expectation value of the kinetic energy operator in the half energy window at left ($[\bar{E}-dE/2,\bar{E}]$) and at right $[\bar{E},\bar{E}+dE/2]$ as a function of $dE$, where $\bar{E}=-1.05$ which has effective temperature $T=3$.
@@
Now we compare our choice of the energy window with the energy window used in [13]. We consider the system size $L=16,N=8$. In the left panel of Figure 11, we plotted the microcanonical expectation values of the kinetic energy operator using two different sizes of the energy window. It appears that $dE=0.4$ and $dE=0.8$ are equally good. The inset of the same figure is the Log-Linear plot of the expectation value in the half energy window at left ($[\bar{E}-dE,\bar{E}]$)and at right $[\bar{E},\bar{E}+dE]$ as a function of $dE$, where $\bar{E}=-1.05$ which has temperature $T=3$. It appears that our choice of $dE$ is a reasonable choice from this plot. But this depends on the choice of the range of $dE$ used. In other word, we believe that choosing a $dE$ from this Log-Linear plot is arbitrary. If we plot a Linear-Linear plot instead of a Log-Linear plot, then we get the right panel of Figure 11. Now, it appears that every choice of $dE$ is equally good (or bad) as long as $dE$ is not very large and comparable to the total size of the energy spectrum.
@@ Appendix D The entropic factor in finite dimensional state space
$\Rightarrow$ $\omega_{0}\sim\frac{4C}{3}\,\frac{1}{\left(2\beta+\frac{1}{C}\frac{\partial C}{\partial T}\right)}$ (42)
@@
$u_{\bar{E}}(\omega)\sim u_{\bar{E}}(\omega_{0})-\frac{(\omega-\omega_{0})^{2}}{2C}\,\left(\beta^{2}_{2}-\beta^{2}_{1}\right)$
@@
Figure 13: Top six subfigures are plots $e^{S(\bar{E}+\omega)-S(\bar{E}+\omega/2)}=e^{u_{\bar{E}}(\omega)}$ as a function of $\omega$ for fixed temperature $T$ ($\bar{E}$ defined in terms of $T$). The different colors are for different system sizes. Bottom two subfigures are plots of Max $e^{u_{\bar{E}}(\omega)}$ as a function of the maxima point $\omega=\omega_{0}$ for different system sizes at fixed temperatures $T=1$ and $T=-1$. The position of the maxima increases with increasing system size, as we expected from the Taylor expansion of $u_{\bar{E}}(\omega)$.
@@ Appendix E Numerical methods and further numerical results
The construction of the Hamiltonian and the other operators are straight-forward in the site basis [37]. After solving for the energy eigenvalues and the energy eigenstates, it is again straight-forward to calculate the matrix elements $\mathcal{O}_{mn}$.
@@
$\mathcal{O}H_{0}\mathcal{O}-\frac{1}{2}\left(\mathcal{O}^{2}H_{0}-H_{0}\mathcal{O}^{2}\right)$ (43)
@@
Panel (a), (b) and (c) in Figure 15 are plots of $f(\beta)$ of the kinetic energy operator for the different system sizes $L=14,N=7$; $L=16,N=8$ and $L=18,N=9$ respectively. The plots support the monotonic behaviour of $f(\beta)$ function. But we can observe something peculiar, the peak of the $\omega=0.1$ plots does not fall on $\beta=0$. Similar departure of the position of the peak from $\beta=0$ is also observed for $\omega=0.5$. We believe that this is a finite size effect. We can see that the peak position, marked by the red tabs on the top $\beta$-axis frames, moves towards $\beta=0$ as the system size increases. This departure of the peak from $\beta=0$ especially in panel (a) suggests that it would be easy to violate the second law of thermodynamics in the smallest system size $L=14,N=7$ by perturbing with the kinetic energy operator with a cut-off of $\omega$ taken to be $\Lambda\sim 0.5$. With such a perturbation, the system would still gain energy starting from a state with effective inverse temperature $\beta=0$. This would be violation of the Kelvin statement. We check if this is the case. Panel (d) is plots of the change in energy as a function of the cut-off $\Lambda$. We considered three initial states - an eigenstate, a linear superposition of eigenstates with energy window $dE=0.1$, and a linear superposition of eigenstates with the microcanonical energy window $dE=0.8$. We find that all of these states generally gain energy from the perturbation. Statistical fluctuations are pronounced due to the small number of dimension of the state space. But we can see that there is a noticeable gain in energy before $\Lambda\sim 0.5$.
Figure 15: Panels (a), (b) and (c) are plots of $f(\beta)$ plots of the kinetic energy operator $\mathcal{O}_{KE}$ with different system sizes $L=14,N=7$ in (a), $L=16,N=8$ in (b), and $L=18,N=9$ in (c). Panel (d) is plots of change in energy $\Delta E$ as a function of $\Lambda$ which is the upper cut-off of $\omega$, starting from states with $\beta=0$, for perturbation with the kinetic energy operator in the system size $L=14,N=7$.
@@
Figure 16: Panel (a) is plots of slope of linear fit of $1/f\partial f/\partial\bar{E}$ for $|\bar{E}|\leq 1$ for the single site spin operator $S^{z}_{L/2}$ with different system sizes $L=14,N=7$ ($7$th site spin operator), $L=16,N=8$ ($8$th site spin operator), and $L=18,N=9$ ($9$th site spin operator). Panels (b), (c) and (d) are plots of the slopes after scaling with different powers of the system size. Panel (c) shows that the slope is decreasing as $L^{-1}$ where $L$ is the system size.

### v12

@@ Appendix C Comparison of canonical and microcanonical ensembles
$S_{\text{can}}=-\sum_{i}p_{i}\log p_{i},\qquad p_{i}=e^{-\beta E_{i}}$ (C.1)
@@
$S_{\mu}=\log W,\qquad\text{$W=$ number of energy eigenstates within $[E_{\beta}-dE/2,E_{\beta}+dE/2]$}$ (C.2)
@@
Figure 10: Left panels: Comparison of canonical entropy $S_{\text{can}}$ and microcanonical entropy $S_{\mu}$ as a function of inverse temperature $\beta$ for different system sizes of chaotic spin chain using different values of the energy window $dE$. Right panels: Plots of $S_{\text{can}}/S_{\mu}$ as a function of inverse temperature $\beta$ for different values of dEdE to study dEdE-dependence of the microcanonical entropy and to find the optimum value of $dE\sim 0.8$.
Figure 11: Finding the value of the energy window dEdE following Rigol_2008nature . These plots are for the system size $L=16,N=8$. Left panel: Plot of microcanonical expectation values of the kinetic energy operator using two different sizes of the energy window $dE=0.4$ and $dE=0.8$. Inset is the Log-Linear plot of the expectation value of the left ($\bar{E}-dE/2,\bar{E}]$) and right $\bar{E},\bar{E}+dE/2]$ averages as a function of $dE$, where $\bar{E}=-1.05$ which has effective temperature $T=3$. Right panel: Linear-Linear plot of the expectation value of the kinetic energy operator in the half energy window at left ($\bar{E}-dE/2,\bar{E}]$) and at right $\bar{E},\bar{E}+dE/2]$ as a function of $dE$, where $\bar{E}=-1.05$ which has effective temperature $T=3$.
@@
Now we compare our choice of the energy window with the energy window used in Rigol_2008nature . We consider the system size $L=16,N=8$. In the left panel of Figure 11, we plotted the microcanonical expectation values of the kinetic energy operator using two different sizes of the energy window. It appears that $dE=0.4$ and $dE=0.8$ are equally good. The inset of the same figure is the Log-Linear plot (following Rigol_2008nature ) of the expectation value in the half energy window at left ($\bar{E}-dE,\bar{E}]$)and at right $\bar{E},\bar{E}+dE]$ as a function of $dE$, where $\bar{E}=-1.05$ which has temperature $T=3$. It appears that our choice of $dE$ is a reasonable choice from this plot. But this depends on the choice of the range of $dE$ used. In other word, we believe that choosing a $dE$ from this Log-Linear plot is arbitrary. If we plot a Linear-Linear plot instead of a Log-Linear plot, then we get the right panel of Figure 11. Now, it appears that every choice of $dE$ is equally good (or bad) as long as $dE$ is not very large and comparable to the total size of the energy spectrum.
@@ Appendix D The entropic factor in finite dimensional state space
$\Rightarrow$ $\omega_{0}\sim\frac{4C}{3}\,\frac{1}{\left(2\beta+\frac{1}{C}\frac{\partial C}{\partial T}\right)}\$(D.1)
@@
$u_{\bar{E}}(\omega)\sim u_{\bar{E}}(\omega_{0})-\frac{(\omega-\omega_{0})^{2}}{2C}\,\left(\beta^{2}_{2}-\beta^{2}_{1}\right)\$
@@
Figure 13: Top six subfigures are plots $e^{S(\bar{E}+\omega)-S(\bar{E}+\omega/2)}=e^{u_{\bar{E}}(\omega)}$ as a function of $\omega$ for fixed temperature $T$ ($\bar{E}$ defined in terms of $T$). The different colors are for different system sizes. Bottom two subfigures are plots of Max $e^{u_{\bar{E}}(\omega)}$ as a function of the maxima point $\omega=\omega_{0}$ for different system sizes at fixed temperatures $T=1$ and T=1T=-1. The position of the maxima increases with increasing system size, as we expected from the Taylor expansion of $u_{\bar{E}}(\omega)$.
@@ Appendix E Numerical methods and further numerical results
The construction of the Hamiltonian and the other operators are straight-forward in the site basis Santos_lecture_2014 . After solving for the energy eigenvalues and the energy eigenstates, it is again straight-forward to calculate the matrix elements $\mathcal{O}_{mn}$.
@@
$\mathcal{O}H_{0}\mathcal{O}-\frac{1}{2}\left(\mathcal{O}^{2}H_{0}-H_{0}\mathcal{O}^{2}\right)$ (E.1)
@@
Panel (a), (b) and (c) in Figure 15 are plots of $f(\beta)$ of the kinetic energy operator for the different system sizes $L=14,N=7$; $L=16,N=8$ and $L=18,N=9$ respectively. The plots support the monotonic behaviour of $f(\beta)$ function. But we can observe something peculiar, the peak of the $\omega=0.1$ plots does not fall on $\beta=0$. Similar departure of the position of the peak from $\beta=0$ is also observed for $\omega=0.5$. We believe that this is a finite size effect. We can see that the peak position, marked by the red tabs on the top $\beta$-axis frames, moves towards $\beta=0$ as the system size increases. This departure of the peak from $\beta=0$ especially in panel (a) suggests that it would be easy to violate the second law of thermodynamics in the smallest system size $L=14,N=7$ by perturbing with the kinetic energy operator with a cut-off of $\omega$ taken to be $\Lambda$. With such a perturbation, the system would still gain energy starting from a state with effective inverse temperature $\beta=0$. This would be violation of the Kelvin statement. We check if this is the case. Panel (d) is plots of the change in energy as a function of the cut-off $\Lambda$. We considered three initial states - an eigenstate, a linear superposition of eigenstates with energy window $dE=0.1$, and a linear superposition of eigenstates with the microcanonical energy window $dE=0.8$. We find that all of these states generally gain energy from the perturbation.
Figure 15: Panels (a), (b) and (c) are plots of $f(\beta)$ plots of the kinetic energy operator $\mathcal{O}_{KE}$ with different system sizes $L=14,N=7$ in (a), $L=16,N=8$ in (b), and $L=18,N=9$ in (c). Panel (d) is plots of change in energy $\Delta E$ as a function of $\Lambda$ which is the upper cut-off of $\omega$, starting from states with $\beta=0$ but different $dE=0,0.1,0.8$, for perturbation with the kinetic energy operator in the system size $L=14,N=7$.
@@
Figure 16: Panel (a) is plots of slope of linear fit of $1/f\partial f/\partial\bar{E}$ for $|\bar{E}|\leq 1$ for the single site spin operator $S^{z}_{L/2}$ with different system sizes $L=14,N=7$ (77th site spin operator), $L=16,N=8$ (88th site spin operator), and $L=18,N=9$ (99th site spin operator). Panels (b), (c) and (d) are plots of the slopes after scaling with different powers of the system size. Panel (c) shows that the slope is decreasing as $L^{-1}$ where $L$ is the system size.